OpenStudy (anonymous):

Can someone please show me the steps to solving (dy/dx)=(4y/x^2) with condition y(-4)=e?

OpenStudy (anonymous):

hiiiiiiiiiiiiiii I can help u

OpenStudy (anonymous):

Okay, great

OpenStudy (holly00d1248):

wheres your work @saynabahmed13

OpenStudy (holly00d1248):

and answer

OpenStudy (anonymous):

lol hold on I am trying to figure it out be

OpenStudy (holly00d1248):

alright LOL

OpenStudy (anonymous):

\[\int dy/4y = \int dx/x^2\]

OpenStudy (anonymous):

Integrate that and finc the constant c using the condition \(y(-4)=e\)

OpenStudy (anonymous):

I am in 6th grade is this even 6th grade math

OpenStudy (anonymous):

I got up to y=4Ce^(-1/x), but I cant find the answer trying to solve for c

OpenStudy (holly00d1248):

@saynabahmed13 this is like 9th 10th grade math

OpenStudy (anonymous):

integrating you get: \[\frac{1}{4}ln(y) = -1/x + c\] \[y = Ce^{\frac{-4}{x}}\]

OpenStudy (anonymous):

Then, \(y(-4)=e\), so \(e = Ce^{-4/-4} \Rightarrow C=1\), so... \(y(x)=e^{-4/x}\)

OpenStudy (anonymous):

I eliminated the ln(y) and had .25y=e^((-1/x)+c) which is when I multiplied and got y=4Ce^(-1/X). That part was not right?

OpenStudy (anonymous):

look, \[\int \frac{dy}{4y} =\frac{1}{4} \int \frac{dy}{y} =\frac{1}{4} ln(y)\]

OpenStudy (anonymous):

u can not do what u did

OpenStudy (anonymous):

first multiply 4 on both sides after apply the exp() in both sides

OpenStudy (anonymous):

Okay, thank you. I understand now.

OpenStudy (anonymous):

No problem ;)